OpenStudy (anonymous):

A gargoyle statue called off the side of a building from a height of 75 m How fast is the statue moving after 1.75 seconds ?

OpenStudy (gajanann):

Free fall Velocity is given by \[V = g*t\] Where, V = velocity of the object. (m/s) g = acceleration due to gravity (m/s/s) t = free fall time. (s) The answer is 9.81*1.75 = 17.17 m/s

OpenStudy (anonymous):

Thank you !!